unI Fald Ines th a veloc 8-fel ㄨㄧㄧㄨㄧㄨ x XxX e the
Answer #1

Part A)

The positive X axis is to right, positive Y axis is upwards and positive Z axis is out of the page.

Unit vectors along Positive X,positive Y and positive Z axes are \hat{i},\hat{j},\hat{k} respectively.

a) In the given figure, magnetic field is directed to right, B= 1.5 i T and velocity vector is making an angle 300 with the field direction, that is 3 * 10(cos 30% + sin 30%) m/s

Magnetic force F=qv × B

F(1.6 1019)3 10(cos 30+sin 30°j) x (1.5i)

F (1.6*1019) 3 105 sin 301.5(-) *3*10sin 30° * 1.

F (3.6 10-4)-k)

Magnitude of magnetic force is F 3.6* 10-14 N and direction of magnetic force is in to the page.

b) Acceleration of proton is a=\frac{F}{m}=\frac{3.6*10^{-14}}{1.67*10^{-27}}=2.16*10^{13}\,m/s^2

-----------------------------

Part B)

In the given figure, magnetic field is directed into the page , B = 1.5(-1) T and velocity vector is to right , that is

2 10 i m/s

1) Path of proton is as shown below.

2) Magnetic force F=qv × B

F= (1.6 * 10-1912 * 10%) × 1.5(-k)

\vec{F}=4.8*10^{-14}\,\hat{j}\,\,N

Magnitude of initial magnetic force on the proton is F-4.8* 10-14 AN and its direction is upwards along positive Y direction.

3) Radius of path followed by proton is r=\frac{mv}{qB}=\frac{1.67*10^{-27}*2*10^5}{1.6*10^{-27}*1.5}=0.00139\,m=0.139\,cm

----------------------------